# iMAD: Intelligent Multi-Agent Debate for Efficient and Accurate LLM Inference

### v1

@@ Abstract
Large Language Model (LLM) agent systems have advanced rapidly, driven by their strong generalization in zero-shot settings. To further enhance reasoning and accuracy on complex tasks, Multi-Agent Debate (MAD) has emerged as a promising framework that engages multiple LLM agents in structured debates to encourage diverse reasoning. However, triggering MAD for every query is inefficient, as it incurs substantial computational (token) cost and may even degrade accuracy by overturning correct single-agent answers. To address these limitations, we propose intelligent Multi-Agent Debate (iMAD), a token-efficient framework that selectively triggers MAD only when it is likely to be beneficial (i.e., correcting an initially wrong answer). To achieve this goal, iMAD learns generalizable model behaviors to make accurate debate decisions. Specifically, iMAD first prompts a single agent to produce a structured self-critique response, from which we extract 41 interpretable linguistic and semantic features capturing hesitation cues. Then, iMAD uses a lightweight debate-decision classifier, trained using our proposed FocusCal loss, to determine whether to trigger MAD, enabling robust debate decisions without test dataset-specific tuning. Through extensive experiments using six (visual) question answering datasets against five competitive baselines, we have shown that iMAD significantly reduces token usage (by up to 92%) while also improving final answer accuracy (by up to 13.5%).
@@ 1 Introduction
Despite this benefit, MAD frameworks face two critical limitations that hinder their practical deployment. First, MAD incurs significantly higher computational costs due to iterative LLM queries–measured in total token usage, which includes both input tokens (prompts to each agent) and output tokens (generated responses). Most MAD systems consume three to five times more tokens than single-agent baselines (liu2024groupdebateenhancingefficiencymultiagent), making them costly to scale (see Insight 3). Second, perhaps more counterintuitively, MAD does not consistently improve response quality. Both prior work (zhang2025stopovervaluingmultiagentdebate) and our own empirical analysis show that in many cases, the single-agent output is already correct, making it redundant to trigger MAD. In other cases, the error in the single-agent response cannot be corrected by MAD, or even worse, triggering MAD may override a correct answer with an incorrect one, resulting in degradation in accuracy (see Insight 3). These observations suggest that while MAD can be beneficial, applying it to every query may be inefficient. Not only does it incur substantial computational costs, but it can also degrade accuracy. Therefore, we need an intelligent and principled mechanism that can selectively trigger MAD only when it is likely to be beneficial. This raises a key question: When should a debate be performed to preserve the benefits of MAD while avoiding unnecessary token costs and potential accuracy degradation?
A straightforward approach is to use the confidence score (typically computed as the average log-probability of the output tokens) to estimate single-agent answer correctness. Leveraging this idea, a recent selective MAD framework, called DOWN, attempts to trigger MAD when the confidence score falls below a threshold (eo2025debatenecessaryadaptivemultiagent). However, it requires a threshold tuned on a subset of the evaluation data, which violates the core zero-shot setting assumption. Moreover, even with a fixed threshold, our empirical analysis reveals that confidence scores alone are not reliable indicators of whether MAD is necessary (see Insight 3). We observe that confidence scores could be high even for incorrect answers, revealing the model’s overconfidence. Also, confidence scores are often misaligned with reasoning uncertainty: responses that contain hesitation cues (such as hedging, contradictions, or shallow reasoning) may still receive a high score. This misalignment causes two undesirable situations: skipping the necessary MAD or triggering it unnecessarily.
@@
(C2) How to design a mechanism that intelligently decides when to trigger MAD for efficient and accurate inference in the zero-shot setting? We identify when MAD is more likely to recover incorrect answers without relying on access to the evaluation dataset. Achieving this requires learning generalizable model behavior, which involves addressing the aforementioned issues of LLM agents: overconfidence in incorrect answers and misalignment between confidence scores and semantic cues of hesitation in the response.
To address these challenges, we propose intelligent Multi-Agent Debate (iMAD), a lightweight debate-triggering framework for token-efficient MAD in the zero-shot setting that selectively triggers MAD only when it is more likely to improve the final answer. The overall workflow of iMAD is illustrated in Fig. 1. We summarize our main contributions along with the key components of iMAD as follows:
@@
  • We evaluate iMAD on three question answering (QA) and three visual question answering (VQA) datasets against five competitive baselines (including two single-agent and three multi-agent frameworks). iMAD reduces token usage by up to 92% while improving accuracy by up to 13.5% through selectively skipping debates that are unnecessary or detrimental. Notably, we train the classifier solely on two representative datasets selected to capture diverse model reasoning behaviors, allowing the classifier to learn generalizable model behavior and perform effectively across six held-out datasets.
@@ 2 Related Work
In a concurrent work, the proposed DOWN framework aims to reduce the token cost of MAD by using LLM-generated confidence scores to determine when to trigger debate (eo2025debatenecessaryadaptivemultiagent). However, selecting an appropriate confidence threshold requires labeled data from the evaluation dataset, which violates the zero-shot setting assumption normally upheld by current single-agent and MAD baselines and limits applicability in real-world scenarios. Moreover, confidence scores alone are not reliable in determining whether triggering a debate will improve the answer (see Insight 3). Although DOWN can reduce token usage by skipping debate on high-confidence responses, it often fails to identify cases where debate is beneficial and misses opportunities to correct recoverable errors.
@@ 3 Key Insights
We quantify the trade-off between accuracy and token costs by comparing CoT (wei2022chain) and MAD (liang-etal-2024-encouraging) across six QA and VQA datasets in Table 1. Consistent with prior findings (liu2024groupdebateenhancingefficiencymultiagent), we observe that MAD achieves a higher accuracy than CoT, with gains ranging from 1.5% (on OKVQA) to 5.3% (on MEDQA). However, MAD consumes three to five times more tokens than CoT, primarily due to routing the same query to multiple agents, each requiring separate input prompts and generating individual responses (chen2024optima; eo2025debatenecessaryadaptivemultiagent). This cost is more pronounced in VQA tasks, where visual inputs further increase token usage. These observations indicate that the accuracy gains come at a high computational cost, making MAD impractical to deploy at scale.
@@
(Insight 3) Confidence scores are not reliable indicators of whether debate is beneficial or not.
We investigate whether LLM-generated confidence scores can effectively indicate when MAD is likely to improve answers (confidence; eo2025debatenecessaryadaptivemultiagent). A natural hypothesis is that answers with a high confidence score may not need MAD; only answers with a low confidence score could benefit from MAD. To test this, we consider confidence scores from various prompting strategies (confidence; eo2025debatenecessaryadaptivemultiagent) and analyze their alignment with cases where MAD corrects initially incorrect single-agent answers. The Cumulative Density Function (CDF) of confidence scores is presented in Fig. 2. We observe that the CDF is highly right-skewed and poorly aligned with answer correctness or debate effectiveness. Notably, incorrect answers often receive high confidence scores, sometimes even exceeding correct answers, indicating a strong bias of overconfidence. Moreover, even many hesitant or shallow responses still receive inflated confidence scores. This misalignment between the confidence score and the uncertainty expressed in reasoning undermines the effectiveness of heuristics that make debate decisions based on confidence scores only.
@@ 4.1 Framework Overview
We extract 41 interpretable linguistic and semantic features from the structured single-agent output, drawing from the question, initial reasoning, and self-critique. These interpretable features capture human-understandable cues to reasoning quality and internal hesitation, including surface-level statistics, readability scores, part-of-speech counts (e.g., nouns, verbs, and adjectives), question-type indicators, and lexical cues of uncertainty, such as hedging and contrast. The feature set enables fine-grained detection of subtle uncertainty cues that are often not reflected in the model’s raw confidence scores. These interpretable features help the classifier identify when MAD is likely to be beneficial by capturing uncertainty cues that are not well aligned with confidence scores. Rather than relying on a subset of features, we combine complementary semantic, syntactic, and pragmatic signals to form a holistic view of model behavior. This is crucial for generalization in zero-shot settings without access to an evaluation dataset. A detailed list of all 41 features is provided in the extended version (see Appendix A.2).
@@
During training, we run the single-agent pipeline on each example to generate an answer and self-critique, and then assign a binary label $y$ where $y=1$ if the answer matches the ground truth and $y=0$ otherwise. The classifier is trained with these binary correctness labels, but its goal is not to replicate these labels. Instead, it aims to identify debatable cases: answers likely to be incorrect but potentially correctable through MAD. This distinction is crucial because confidently incorrect answers, often resulting from coherent but flawed reasoning, are unlikely recoverable through MAD.
During inference, we apply a decision threshold $\tau\in(0,1)$, tuned on a validation set, to the predicted score $p$: if $p<\tau$, MAD is triggered; otherwise, the original single-agent answer is retained. A high $p$ indicates triggering MAD is unnecessary, while a low $p$ flags uncertain and potentially recoverable errors that warrant MAD. This selective mechanism enables iMAD to devote computational resources to the most uncertain and error-prone cases, reducing token costs while improving accuracy. The classifier is trained using our proposed FocusCal loss, detail in Section 4.2.
@@ 4.2 Debate Decision Classifier with FocusCal Loss
This component places the strongest penalty on the cases where the classifier assigns a high predicted score $p$ to an answer that is factually incorrect (i.e., labeled $y=0$). These are exactly the instances where MAD should be triggered but was mistakenly skipped. To address this, we adopt an asymmetric focal loss defined as
$L_{\text{AF}}(y,p):=\begin{cases}-\alpha_{1}(1-p)^{\gamma}\log(p),&\text{if }y=1;\\ -\alpha_{0}~p^{\gamma}\log(1-p),&\text{if }y=0,\end{cases}$ (6)
@@
To further align the model’s predicted score $p$ with its uncertainty score $u$, we introduce a regularization loss term that penalizes misalignment between them. This uncertainty score $u$ reflects the model’s internal hesitation: a high value of $u$ indicates distributed uncertainty across semantic interpretations, while a low value of $u$ indicates peaked, confident predictions. Although $p$ expresses the overall correctness belief, $u$ offers a complementary perspective by capturing the uncertainty or hesitation embedded in the response features. To reconcile these two signals, we define Confidence Penalty as
@@
This confidence penalty term penalizes underconfidence for correct answers (when high uncertainty occurs with a low $p$) and overconfidence for incorrect ones (when low uncertainty occurs with a high $p$). By enforcing the consistency between the predicted score $p$ and the uncertainty score $u$, this loss mitigates the misalignment between the predicted score $p$ and the semantic hesitation signals, leading to more reliable debate triggering decisions.
@@
Table 3: Accuracy (Acc) and average token cost (# Token) per question comparison of iMAD and baselines across datasets.
@@
To improve the reliability of $p$ as a debate-triggering score, we incorporate ECE to encourage its alignment with the empirical correctness. Suppose the dataset contains $N$ instances indexed by $i\in\{1,\dots,N\}$, where each instance $i$ has a predicted score $p^{(i)}\in[0,1]$ and a binary ground-truth label $y^{(i)}\in\{0,1\}$. Based on $p^{(i)}$, We divide the interval $[0,1]$ into $B$ equal-width bins and assign each instance $i$ to a bin $b\in\{1,\dots,B\}$. Let $\mathcal{I}_{b}$ denote the set of indices $i$ whose predicted scores falls into bin $b$. We then compute ECE as:
@@ 5 Evaluation
In this section, we evaluate iMAD across six datasets, measuring both token efficiency (i.e., token usage per question) and accuracy (i.e., final answer correctness). We further analyze the quality of iMAD’s debate decisions on whether to trigger or skip MAD, and how often these decisions lead to beneficial outcomes. Finally, we present ablation studies on the impact of structured self-critique prompting and FocusCal loss in the extended version (see Appendices C.1 and  C.2, respectively).
5.1 Experimental Setup
Datasets. We conduct experiments on six diverse datasets spanning both textual QA and image-text VQA domains. The QA benchmarks include (i) MedQA: USMLE-style medical QA (jin2020disease); (ii) MMLU: professional exam questions (hendrycks2020measuring); and (iii) GSM8K: grade-school math word problems (cobbe2021gsm8k). For VQA, we use (iv) OKVQA: emphasizes visual questions that require external knowledge (okvqa); (v) VQA-v2: natural image QA with reduced bias (balanced_vqa_v2); and (vi) ScienceQA: multimodal science questions from school curricula (lu2022learn).
@@ 5.1 Experimental Setup
We compare iMAD with five strong baselines across single-agent, full-debate MAD, and selective MAD approaches. For single-agent methods, we include: (i) CoT (wei2022chain); and (ii) SC (wang2022self), which runs CoT five times and selects the answer via majority voting. For full-debate MAD approaches that trigger debate for all instances, we consider: (iii) MAD (liang-etal-2024-encouraging), using agents with distinct personas in a three-agent setup; and (iv) GD (liu2024groupdebateenhancingefficiencymultiagent), which clusters 5 agents into subgroups for parallel discussion, followed by 3 rounds of inter-group consensus voting. For selective MAD triggering, we evaluate: (v) DOWN (eo2025debatenecessaryadaptivemultiagent). Unlike other baselines, DOWN requires labeled evaluation data to tune its threshold. For fair comparison, we use a threshold of 0.8 as reported most effective in the original paper.
@@
We use a lightweight MLP with six fully connected layers of 200 hidden units each, the batch normalization, ReLU activations, and a dropout rate of 0.2. We train the model for 50 epochs using the Adam optimizer (adam) with learning rate of 0.001 on standardized features via StandardScaler (de2023choice). To support task-agnostic generalization, we train the classifier on PubMedQA (pubmedqa) and GQA (GQA), which are not included in the evaluation. We set the FocusCal loss hyperparameters as $\alpha_{\text{0}}=2.0$, $\alpha_{\text{1}}=1.0$, $\gamma=2$, $\lambda=6$, $\mu=5$, and $B=15$ bins for ECE. The debate threshold is set to $\tau=0.7$. We chose these hyperparameters via grid search on a held-out validation set to ensure stable training and strong performance across all datasets.
@@
For LLM-based prompting, we use Gemini 2.0 Flash as the LLM agents for the primary results. We also evaluate GPT-5 nano and Qwen 3.0 (see Appendix C.3 for details and results). By default, we use a temperature of 0.0 to ensure deterministic outputs and a maximum sequence length of 512 tokens. We present the prompt templates details in the extended version (see Appendix A.1). We conducted all MLP training and inference using a single NVIDIA RTX 4090 GPU.
@@
Table 4: Breakdown of iMAD debate decisions by whether MAD was skipped or triggered, and whether the decision was beneficial (Good) or harmful (Bad) based on changes in accuracy or token cost across datasets.

### v2

@@ Abstract
Large Language Model (LLM) agent systems have advanced rapidly, driven by their strong generalization in zero-shot settings. To further enhance reasoning and accuracy on complex tasks, Multi-Agent Debate (MAD) has emerged as a promising framework that engages multiple LLM agents in structured debates to encourage diverse reasoning. However, triggering MAD for every query is inefficient, as it incurs substantial computational (token) cost and may even degrade accuracy by overturning correct answers from single-agent. To address these limitations, we propose intelligent Multi-Agent Debate (iMAD), a token-efficient framework that selectively triggers MAD only when it is likely to be beneficial (i.e., correcting an initially wrong answer). To achieve this goal, iMAD learns generalizable model behaviors to make accurate debate decisions. Specifically, iMAD first prompts a single agent to produce a structured self-critique response, from which we extract 41 interpretable linguistic and semantic features capturing hesitation cues. Then, iMAD uses a lightweight debate decision classifier, trained using our proposed FocusCal loss without test-dataset-specific tuning, to make robust zero-shot debate decisions. Through extensive experiments using six (visual) question answering datasets against five competitive baselines, we show that iMAD significantly reduces token usage (by up to 92%) while also improving final answer accuracy (by up to 13.5%).
Codehttps://github.com/Fanwei100/iMAD
@@ 1 Introduction
Despite this benefit, MAD frameworks face two critical limitations that hinder their practical deployment. First, due to iterative LLM queries, MAD incurs significantly higher computational costs–measured in total token usage, which includes both input tokens (prompts to each agent) and output tokens (generated responses). Most MAD systems consume three to five times more tokens than single-agent baselines (liu2024groupdebateenhancingefficiencymultiagent), making them costly to scale (see Insight 3). Second, perhaps more counterintuitively, MAD does not consistently improve response quality. Both prior work (zhang2025stopovervaluingmultiagentdebate) and our own empirical analysis show that in many cases, the single-agent output is already correct, making it redundant to trigger MAD. In other cases, the error in the single-agent response cannot be corrected by MAD, or even worse, triggering MAD may override a correct answer with an incorrect one, resulting in degraded accuracy (see Insight 3). These observations suggest that while MAD can be beneficial, applying it to every query may be inefficient. Not only does it incur substantial computational costs, but it can also degrade accuracy. Therefore, we need an intelligent and principled mechanism that can selectively trigger MAD only when it is likely to be beneficial. This raises a key question: When should a debate be performed to preserve the benefits of MAD while avoiding unnecessary token costs and potential accuracy degradation?
A straightforward approach is to use the confidence score (typically computed as the average log-probability of the output tokens) to estimate single-agent answer correctness. Leveraging this idea, a concurrent work recently proposed a selective MAD framework, called DOWN, which attempts to trigger MAD when the confidence score falls below a threshold (eo2025debatenecessaryadaptivemultiagent). However, it requires a threshold tuned on a subset of the evaluation data, which violates the core zero-shot setting assumption. Moreover, even with a fixed threshold, our empirical analysis reveals that confidence scores alone are not reliable indicators of whether MAD is necessary (see Insight 3). We observe that confidence scores could be high even for incorrect answers, revealing the model’s overconfidence. Also, confidence scores are often misaligned with reasoning uncertainty: responses that contain hesitation cues (such as hedging, contradictions, or shallow reasoning) may still receive a high score. This misalignment causes two undesirable situations: skipping the necessary MAD or triggering it unnecessarily.
@@
(C2) How to design a mechanism that intelligently decides when to trigger MAD for efficient and accurate inference in the zero-shot setting? One needs to identify when MAD is more likely to recover incorrect answers without relying on the evaluation dataset. This requires learning generalizable model behaviors, which involves addressing the aforementioned issues of LLM agents: overconfidence in incorrect answers and misalignment between confidence scores and semantic cues of hesitation in the response.
To address these challenges, we propose intelligent Multi-Agent Debate (iMAD), a lightweight debate-triggering framework for token-efficient MAD in the zero-shot setting. iMAD selectively triggers MAD only when it is more likely to improve the final answer. The overall workflow of iMAD is illustrated in Fig. 1. We summarize our main contributions along with the key components of iMAD as follows:
@@
  • We evaluate iMAD on three question answering (QA) and three visual question answering (VQA) datasets against five competitive baselines (including two single-agent and three multi-agent frameworks). iMAD reduces token usage by up to 92% while improving accuracy by up to 13.5% through selectively skipping debates that are unnecessary or detrimental. Notably, we train the classifier solely on two representative datasets selected to capture diverse model reasoning behaviors, allowing the classifier to learn generalizable model behaviors and perform effectively across six held-out datasets.
@@ 2 Related Work
The concurrent work DOWN aims to reduce MAD token costs by using LLM-generated confidence scores to decide when to trigger debate (eo2025debatenecessaryadaptivemultiagent). However, selecting an appropriate confidence threshold requires labeled evaluation data, which violates the zero-shot setting assumption normally upheld by current single-agent and MAD baselines and limits real-world applicability. Moreover, confidence scores alone are unreliable in determining whether triggering a debate will improve the answer (see Insight 3). Although DOWN saves tokens by skipping debate on high-confidence responses, it often fails to identify cases where debate is beneficial and misses opportunities to correct recoverable errors.
@@ 3 Key Insights
We quantify the trade-off between accuracy and token costs by comparing CoT (wei2022chain) and MAD (liang-etal-2024-encouraging) across six QA and VQA datasets in Table 1. Consistent with prior findings (liu2024groupdebateenhancingefficiencymultiagent), we observe that MAD achieves a higher accuracy than CoT, with gains ranging from 1.5% (on OKVQA) to 5.3% (on MEDQA). However, MAD consumes 3–5 times more tokens than CoT, mainly due to routing the same query to multiple agents, each requiring separate input prompts and generating individual responses (chen2024optima; eo2025debatenecessaryadaptivemultiagent). This cost is more pronounced in VQA tasks, where visual inputs further increase token usage. These observations indicate that the accuracy gains come at a high token cost, making MAD impractical to deploy at scale.
@@
(Insight 3) Confidence scores are unreliable indicators of whether debate is beneficial or not.
We investigate whether LLM-generated confidence scores can effectively indicate when MAD is likely to improve answers (confidence; eo2025debatenecessaryadaptivemultiagent). A natural hypothesis is that answers with a high confidence score may not need MAD; only answers with a low confidence score could benefit from MAD. To test this, we consider confidence scores from various prompting strategies (confidence; eo2025debatenecessaryadaptivemultiagent) and analyze their alignment with cases where MAD corrects initially incorrect single-agent answers. In Fig. 2, we observe that the Cumulative Density Function (CDF) of confidence scores is highly right-skewed and poorly aligned with answer correctness or debate effectiveness. Notably, incorrect answers often receive high confidence scores, sometimes even exceeding correct answers, indicating a strong bias of overconfidence. Moreover, even many hesitant or shallow responses still receive inflated confidence scores. This misalignment between the confidence score and the uncertainty in reasoning undermines the effectiveness of heuristics that make debate decisions based on confidence scores only.
@@ 4.1 Framework Overview
We extract 41 interpretable linguistic and semantic features from the structured single-agent output, drawing from the question, initial reasoning, and self-critique. These interpretable features capture human-understandable cues to reasoning quality and internal hesitation, including surface-level statistics, readability scores, part-of-speech counts (e.g., nouns, verbs, and adjectives), question-type indicators, and lexical cues of uncertainty, such as hedging and contrast. The feature set enables fine-grained detection of subtle uncertainty cues that are often not reflected in the model’s raw confidence scores. These interpretable features help the classifier identify when MAD is likely to be beneficial by capturing uncertainty cues that are not well aligned with confidence scores. Rather than relying on a subset of features, we combine complementary semantic, syntactic, and pragmatic signals to form a holistic view of model behavior. This is crucial for generalization in zero-shot settings without access to an evaluation dataset. A detailed list of all 41 features is provided in Appendix A.2.
@@
During training, we run the single-agent pipeline on each instance to generate an answer and self-critique, and then assign a binary label $y$ where $y=1$ if the answer matches the ground truth and $y=0$ otherwise. The classifier is trained with these binary correctness labels, but its goal is not to replicate these labels. Instead, it aims to identify debatable cases: answers likely to be wrong but potentially correctable through MAD. This distinction is crucial because confidently incorrect answers, often resulting from coherent but flawed reasoning, are unlikely recoverable via MAD.
During inference, we apply a decision threshold $\tau\in(0,1)$, tuned on a validation set, to the predicted score $p$: if $p<\tau$, MAD is triggered; otherwise, the original single-agent answer is retained. A high $p$ indicates triggering MAD is unnecessary, while a low $p$ flags uncertain and potentially recoverable errors that warrant MAD. This selective mechanism enables iMAD to devote computational resources to the most uncertain and error-prone cases, reducing token costs while improving accuracy. The classifier is trained using our proposed FocusCal loss, detailed in Section 4.2.
@@ 4.2 Debate Decision Classifier with FocusCal Loss
This term places the strongest penalty on the cases where the classifier assigns a high predicted score $p$ to an answer that is factually incorrect (i.e., labeled $y=0$). These are exactly the instances where MAD should be triggered but was mistakenly skipped. To address this, we adopt an asymmetric focal loss defined as:
$L_{\text{AF}}(y,p):=\begin{cases}-\alpha_{1}(1-p)^{\gamma}\log(p),&\text{if }y=1,\\ -\alpha_{0}~p^{\gamma}\log(1-p),&\text{if }y=0,\end{cases}$ (6)
@@
To further align the model’s predicted score $p$ with its uncertainty score $u$, we introduce a regularization loss term that penalizes misalignment between them. This uncertainty score $u$ reflects the model’s internal hesitation: a high value of $u$ indicates distributed uncertainty across semantic interpretations, while a low value of $u$ indicates peaked, confident predictions. Although $p$ expresses the overall correctness belief, $u$ offers a complementary perspective by capturing the uncertainty or hesitation embedded in the response features. To reconcile these two signals, we define Confidence Penalty as follows:
@@
This confidence penalty term penalizes under-confidence for correct answers (high uncertainty occurs with a low $p$) and overconfidence for incorrect ones (low uncertainty occurs with a high $p$). By enforcing consistency between the predicted score $p$ and the uncertainty score $u$, this loss mitigates the misalignment between $p$ and the semantic hesitation signals, leading to more reliable debate triggering decisions.
@@
Table 3: Accuracy (Acc) and average token cost (# Token) per question comparison of iMAD and baselines across datasets. Bold values indicate the best result in each row, and underlined values indicate the second best.
@@
To improve the reliability of $p$ as a debate-triggering score, we incorporate ECE to encourage its alignment with empirical correctness. Suppose the dataset contains $N$ instances indexed by $i\in\{1,\dots,N\}$, where each instance $i$ has a predicted score $p^{(i)}\in[0,1]$ and a binary ground-truth label $y^{(i)}\in\{0,1\}$. Based on $p^{(i)}$, We divide the interval $[0,1]$ into $B$ equal-width bins and assign each instance $i$ to a bin $b\in\{1,\dots,B\}$. Let $\mathcal{I}_{b}$ denote the set of indices $i$ whose predicted scores fall into bin $b$. We then compute ECE as:
@@ 5 Evaluation
We report token efficiency (i.e., token usage per question) and accuracy (i.e., final answer correctness) across six datasets. We also analyze iMAD’s debate decisions on whether to trigger or skip MAD, and how often these decisions lead to beneficial outcomes. We further provide additional analysis of token efficiency and inference time in Appendix B, and present additional results, including ablation studies on structured self-critique prompting and FocusCal loss, as well as cross-LLM evaluation, in Appendix C.
Datasets.
We evaluate on six datasets across textual QA and image-text VQA. The QA benchmarks include (i) MedQA: USMLE-style medical QA (jin2020disease); (ii) MMLU: professional exam questions (hendrycks2020measuring); and (iii) GSM8K: grade-school math word problems (cobbe2021gsm8k). For VQA, we use (iv) OKVQA: emphasizes visual questions that require external knowledge (okvqa); (v) VQA-v2: natural image QA with reduced bias (balanced_vqa_v2); and (vi) ScienceQA: science questions from school curricula (lu2022learn).
@@
We compare iMAD with five strong baselines across single-agent, full-debate MAD, and selective MAD approaches. For single-agent methods, we include: (i) CoT (wei2022chain); and (ii) SC (wang2022self), which runs CoT five times and selects the answer via majority voting. For full-debate MAD approaches that trigger debate for all instances, we consider: (iii) MAD (liang-etal-2024-encouraging), using agents with distinct personas in a three-agent setup; and (iv) GD (liu2024groupdebateenhancingefficiencymultiagent), which clusters 5 agents into subgroups for parallel discussion, followed by 3 rounds of inter-group consensus voting. For selective MAD triggering, we evaluate: (v) DOWN (eo2025debatenecessaryadaptivemultiagent). Unlike other baselines, DOWN requires labeled evaluation data to tune its threshold. For fair comparison, we use a threshold of 0.8 as reported to be most effective in the original paper.
@@
We use a lightweight MLP with six fully connected layers of 200 hidden units each, batch normalization, ReLU activations, and a dropout rate of 0.2. We train the classifier for 50 epochs using the Adam optimizer (adam) with a learning rate of 0.001 on standardized features via StandardScaler (de2023choice). To support task-agnostic generalization, we use PubMedQA (pubmedqa) and GQA (GQA) datasets for training, which are not included in the evaluation. We set the FocusCal loss hyperparameters as $\alpha_{\text{0}}=2.0$, $\alpha_{\text{1}}=1.0$, $\gamma=2$, $\lambda=6$, $\mu=5$, and $B=15$ bins for ECE. The debate threshold is set to $\tau=0.7$. We chose these hyperparameters via grid search on a held-out validation set to ensure stable training and strong performance across all datasets.
@@
For LLM-based prompting, we use Gemini 2.0 Flash as the LLM agents for the primary results and also use GPT-5 nano and Qwen 3.0 (see Appendix C.3 for details and results). By default, we use a temperature of 0.0 to ensure deterministic outputs and a maximum of 512 tokens. We present the detailed prompt templates in Appendix A.1. We conducted MLP training and inference using a single NVIDIA RTX 4090 GPU.
@@
Table 4: Breakdown of iMAD debate decisions by whether MAD was skipped or triggered, and whether the decision was beneficial (Good) or harmful (Bad).